Juila12001:

does number 4 equal 16 or 17

Juila12001:

https://ibb.co/cr6Ez7

iGreen:

Cosine of the angle is equal to Adjacent/Hypotenuse, so it should be: \(\sf cos(48^o)=\dfrac{12}{x}\)

iGreen:

Now solve for 'x'.

Juila12001:

don't still get 8.0292

iGreen:

What? \(\sf x=\dfrac{12}{cos(48^o)}\)

iGreen:

What do you get when you simplify the right-hand side?

Juila12001:

don't find the degree of 48 which is 0.6691 and multiple that with 12

Juila12001:

so is my answer right?